What minimum heat is needed to bring 250g of water at 50?C to the boiling point and completely boil it away? The specific heat of water is 4190 J/(kg?K) and its heat of vaporization is 22.6
Answer #1

from 50 to 100
= m c dT = .25*50*4186 = 52325 j

for boil out = Lv . m
= 22.6*10^5 *0.25
=565000

total = sum of both = 6.17*10^5 j